### v1

# PSLM: Parallel Generation of Text and Speech with LLMs for Low-Latency Spoken Dialogue Systems

Abstract
Multimodal language models that process both text and speech have a potential for applications in spoken dialogue systems. However, current models face two major challenges in response generation latency: (1) generating a spoken response requires the prior generation of a written response, and (2) speech sequences are significantly longer than text sequences. This study addresses these issues by extending the input and output sequences of the language model to support the parallel generation of text and speech. Our experiments on spoken question answering tasks demonstrate that our approach improves latency while maintaining the quality of response content. Additionally, we show that latency can be further reduced by generating speech in multiple sequences. Demo samples are available at https://rinnakk.github.io/research/publications/PSLM.
PSLM: Parallel Generation of Text and Speech with LLMs
for Low-Latency Spoken Dialogue Systems
Kentaro Mitsui, Koh Mitsuda, Toshiaki Wakatsuki, Yukiya Hono, Kei Sawada rinna Co., Ltd., Tokyo, Japan {kemits,kohmi,towaka,yuhono,keisawada}@rinna.co.jp
1 Introduction
Spoken dialogue systems have been developed for many years to achieve natural human-computer interaction (McTear, 2002; Jokinen and McTear, 2009; Chen et al., 2017). Traditionally, these systems consist of several components: Automatic Speech Recognition (ASR), Response Generation (RG), and Text-to-Speech (TTS). Various methods for RG have been proposed with the advancements in Large Language Models (LLMs) (Wang et al., 2023a; Yi et al., 2024). More recently, the application of LLMs to ASR (e.g., Wang et al. 2023b; Hono et al. 2024; Fathullah et al. 2024) and TTS (Wang et al., 2023b; Hao et al., 2023) has attracted much attention, leading to the development of multimodal LLMs capable of end-to-end spoken language communication (Zhang et al., 2023; Nachmani et al., 2024).
Zhang et al. (2023) proposed SpeechGPT, an LLM that receives speech questions (SQ) as speech tokens, which are discrete representations extracted from raw waveforms, and sequentially generates text questions (TQ), text answers (TA), and speech answers (SA). Figure 1 (a) illustrates their approach called Chain-of-Modality (CoM) prompting. Spectron (Nachmani et al., 2024) follows this prompting style but directly handles speech spectrograms. Although these methods can generate high-quality responses, they face two major challenges in terms of response latency. First, generating SA requires the prior generation of TQ and TA. Second, speech sequences are much longer than text sequences [^1] Actual sequence lengths are provided in Appendix A..
Figure 1: (a) Chain-of-Modality prompting necessitates generating text questions (TQ) and text answers (TA) from speech questions (SQ) before producing speech answers (SA). (b) Our Parallel Speech Language Model (PSLM) enables the parallel decoding of TA and SA, reducing overall latency. (c) Introducing multiple speech streams further accelerates the generation of SA.
In this study, we propose Parallel Speech Language Model (PSLM), an LLM with multiple input-output sequences to handle both text and speech tokens, enabling their parallel generation. To emphasize their parallel processing capabilities, we will refer to these sequences as “streams”. As described in Figure 1 (b), PSLM begins to generate SA immediately after the end of SQ tokens, which can reduce overall latency. This leads to our first research question (RQ1): Can PSLM improve latency while maintaining the response quality achieved by CoM prompting? Additionally, we address the second challenge by introducing multiple speech streams to decode multiple speech tokens in a single step, as described in Figure 1 (c). This brings us to the second research question (RQ2): Do multiple speech streams sacrifice the response quality? Addressing these questions will pave the way for more advanced and responsive applications of spoken dialogue systems.
2 PSLM
2.1 Speech Discretization
Speech Tokenization
Extracting discrete speech tokens from raw waveforms enables language models to handle speech in the same manner as text tokens. Self-supervised learning has been widely used for speech tokenization due to its ability to extract spoken content from raw waveforms (e.g., Rubenstein et al. 2023; Chou et al. 2023; Hassid et al. 2023). Following Zhang et al. (2023), we employ Hidden-Unit BERT (HuBERT) (Hsu et al., 2021) for speech tokenization.
Speech Detokenization
In contrast to text tokenization, which is uniquely recoverable, speech tokenization largely discards the information of raw waveforms. Two major approaches have been proposed to solve this problem. The first approach uses a neural vocoder for directly reconstructing raw waveforms from speech tokens (e.g., Zhang et al. 2023; Chou et al. 2023; Hassid et al. 2023). The second approach uses a pretrained neural audio codec, which requires an additional module to predict the codec’s tokens (e.g., Rubenstein et al. 2023; Zhang et al. 2024). We adopt the first approach to reduce overall latency using HiFi-GAN (Kong et al., 2020), a non-autoregressive neural vocoder that efficiently generates high-fidelity waveforms.
2.2 Integrating LMs with a Speech Stream
PSLM is built on top of a pretrained decoder-only Transformer (Vaswani et al., 2017). We add new input embedding and output projection layers to process speech tokens, while the structure of the intermediate Transformer layers remains unchanged. We randomly initialize the weights of new embedding and projection layers.
A challenge of joint text-speech modeling lies in the mismatch in their lengths. In this study, we simply right-pad TQ and TA sequences with special [TEXT-PAD] tokens to align their lengths with those of the SQ and SA sequences, respectively. With this approach, we force the PSLM to learn alignment between text and speech. Our PSLM is trained by minimizing the sum of cross entropy losses for each stream. We include prompt tokens, comprising TQ and SQ, in the loss calculation. During inference, PSLM receives these prompt tokens and generates TA and SA in parallel.
2.3 Introducing Multiple Speech Streams
For further acceleration, we introduce multiple speech streams to PSLM. Assume that PSLM has $1+S$ streams, one for text tokens and $S$ for speech tokens. We first reshape the original speech token sequence of length $N$ into a matrix with dimensions $(N/S,S)$, and use the $s$-th column as the $s$-th speech stream ($s\in\{1,\ldots,S\})$. This reduces the sequence length to $1/S$, which leads to approximately an $S$ times speedup. During training, simply summing the cross entropy losses for each stream makes the loss of text tokens less dominant, leading to poor text generation quality. Therefore, we introduce a weighted loss, where we multiply the loss for speech streams by $1/S$ to balance the weight of losses for text and speech streams.
2.4 Streaming Inference with HiFi-GAN
Following Chen et al. (2022), we use HiFi-GAN for streaming inference; specifically, we provide partial speech tokens to generate waveform fragments. In this study, we use non-causal convolution to maintain high speech quality. Therefore, the first speech fragment can be generated once $N_{\textrm{offset}}=\lfloor R/2\rfloor+1$ tokens are decoded, where $R$ denotes the receptive field of HiFi-GAN. Implementation details can be found in Appendix 2.
2.5 Overall Latency
We define latency as the delay between the end of the user’s utterance and the system’s initial response. The latency of conventional CoM-based systems $L_{\textrm{CoM}}$ can be represented as follows:
$L_{\textrm{CoM}}$ $=D_{\textrm{s2t}}+D_{\textrm{SQ}}+\frac{N_{\textrm{dec}}}{P}+D_{\textrm{t2s}}$ (1)
$N_{\textrm{dec}}$ $=N_{\textrm{TQ}}+N_{\textrm{TA}}+N_{\textrm{offset}}$ (2)
where $D_{\textrm{s2t}}$, $D_{\textrm{SQ}}$, and $D_{\textrm{t2s}}$ denote the delays of speech tokenization, the prefill phase in LMs, and speech detokenization, respectively; $N_{\textrm{TQ}}$ and $N_{\textrm{TA}}$ denote the number of tokens in TQ and TA, respectively; and $P$ denotes the tokens per second (TPS) during the decode phase in LMs.
Our PSLM eliminates the need for generating TQ and TA beforehand, although it requires to run external ASR to obtain TQ. Hence, its latency $L_{\textrm{PSLM}}$ can be represented as follows:
$L_{\textrm{PSLM}}=D_{\textrm{ASR}}+D_{\textrm{SQ}}+\frac{N_{\textrm{offset}}}{P\cdot S}+D_{\textrm{t2s}}$ (3)
where $D_{\textrm{ASR}}$ denotes the ASR delay. Here $D_{\textrm{s2t}}$ is omitted because speech tokenization can be performed in parallel with ASR.
3 Experimental Setup
3.1 Dataset
We used an internal dataset comprising 1.8M written QA pairs for training and a public Japanese instruction dataset (Hayashibe, 2023) comprising 669 written QA pairs for evaluation. We filtered out samples with more than 140 characters in TA, obtaining 396 samples for evaluation. For both sets, we constructed a spoken question answering (SQA) dataset by synthesizing SQ and SA using a well-trained single-speaker TTS system based on VITS (Kim et al., 2021).
3.2 Configuration
Tokenization and Detokenization
For text tokenization, we used the tokenizer with a vocabulary size of 151,936 from rinna/nekomata-7b [^2] https://huggingface.co/rinna/nekomata-7b. For speech tokenization, we applied $k$-means clustering with $k=512$ to 12-th layer features from rinna/japanese-hubert-base [^3] https://huggingface.co/rinna/japanese-hubert-base (Sawada et al., 2024), obtaining 50 speech tokens per second. For speech detokenization, we trained discrete unit-based HiFi-GAN (Polyak et al., 2021) using pairs of synthesized speech waveforms of SQ and SA and their corresponding speech tokens. For ASR, Whisper large-v3 (Radford et al., 2023) with faster-whisper [^4] https://github.com/SYSTRAN/faster-whisper was used throughout our experiments.
Language Modeling
We used rinna/nekomata-7b, a 32-layer 4096-hidden-size Transformer LM that was continuously pretrained from Qwen-7B (Bai et al., 2023) on Japanese text, as the backbone of our models. We implemented our models using the GPT-NeoX library (Andonian et al., 2023). Unless otherwise noted, models were trained for 50k steps with a batch size of 16 on 8 NVIDIA A100 GPUs using an Adam optimizer (Kingma and Ba, 2015) with a peak learning rate set to 1e-5. During inference, we set the temperature to 0.8 and applied top-$k$ and top-$p$ sampling with $k=60$ and $p=0.8$.
3.3 Baselines
We involved three CoM-based baselines, which share the model weights but differ in their prompts during decoding: (1) CoM-SQ receives only SQ, (2) CoM-ASR receives SQ and transcribed TQ, and (3) CoM receives SQ and gold TQ.
3.4 Evaluation Metrics
ChatGPT Scores
We used OpenAI’s GPT-3.5 Turbo API to evaluate response quality on a 5-point scale from 1 (bad) to 5 (excellent). The prompt is described in Appendix C. We report the scores for TA and the transcription of SA as T-score and S-score, respectively.
Character Error Rate (CER)
We calculated the character error rate between the generated TA and the transcription of SA to assess their alignment.
Failure Rate (FR)
We counted failure cases such as (1) no [EOS] token was generated before the total sequence length reached 2048, or (2) tokens were generated in the wrong modality, i.e., speech tokens in TQ and TA, or text tokens in SA.
Latency
We simulated latency according to Equations 2 and 3 for each sample in the evaluation set, and reported the median values. Further details can be found in Appendix E.1.
Table 1: Automatic evaluation results. T-score and S-score represent the ChatGPT-based score for TA and transcribed SA, respectively. FR denotes the failure rate. Latency values in parentheses represent inputs involving gold TQ.
[…]
Ground Truth 4.00$\pm$0.02 3.58$\pm$0.06 7.35
CoM SQ $\rightarrow$ TQ (Gold) TA $\rightarrow$ SA 3.50$\pm$0.09 3.27$\pm$0.09 12.12 6.28 (0.67)
PSLM SQ, TQ (Gold) TA, SA 3.50$\pm$0.08 3.22$\pm$0.09 5.05 5.25 (0.34)
CoM-SQ SQ TQ $\rightarrow$ TA $\rightarrow$ SA 3.12$\pm$0.11 2.94$\pm$0.10 15.91 7.83 1.03
CoM-ASR SQ $\rightarrow$ TQ (ASR) TA $\rightarrow$ SA 3.27$\pm$0.10 3.07$\pm$0.09 13.13 6.18 0.92
PSLM-ASR SQ, TQ (ASR) TA, SA 3.34$\pm$0.09 3.05$\pm$0.10 6.31 6.05 0.54
PSLM-2x SQ, TQ (Gold) TA, SA 3.50$\pm$0.08 3.20$\pm$0.09 4.29 6.39 (0.20)
PSLM-3x SQ, TQ (Gold) TA, SA 3.28$\pm$0.10 2.99$\pm$0.10 7.07 6.09 (0.15)
Human Rating
We also conducted two subjective evaluations: one for text and the other for speech. In the text evaluation, we presented pairs of gold TQ and generated TA, along with the criteria used in ChatGPT-based evaluation. In the speech evaluation, we presented gold SQ and generated SA successively, along with their TQ and TA, and asked the raters to evaluate (1) how natural the SA is as the speech of the TA, and (2) whether the response is fast enough. The duration of silence between SQ and SA was simulated in the manner described in Section 2.5, except for the Ground Truth where the silence duration was set to 200ms, the average turn-taking gap in human conversation (Levinson and Torreira, 2015). Scores were rated on a 5-point scale. Fifty samples were randomly chosen from the evaluation set, and twenty in-house workers rated twenty samples each.
4 Results and Discussion
4.1 Automatic Evaluation
Comparison with Baselines
To answer RQ1, we compared the proposed method in two conditions, PSLM and PSLM-ASR, with the baselines described in Section 3.3. PSLM receives SQ and gold TQ, while PSLM-ASR receives SQ and transcribed TQ. Table 1 summarizes the results. When gold TQ was given, PSLM achieved comparable scores to CoM and significantly improved latency. A similar trend was observed under more practical conditions where gold TQ was not available (PSLM-ASR vs. CoM-ASR). However, their scores were lower than those with gold TQ, and CoM-SQ faced greater degradation. These results suggest that ASR performance is crucial for response quality, and CoM-SQ seems to have produced more ASR errors than Whisper. Nevertheless, we conclude that PSLM maintains the response quality of CoM (RQ1). We also found that PSLM-based methods achieved lower FRs than CoM-based ones. Each stream of PSLM is dedicated to a single modality, which could have reduced the failures in generation. Furthermore, methods other than CoM-SQ marked lower CERs than Ground Truth. From this result, we confirmed that both CoM and PSLM can generate appropriate SA corresponding to TA.
Multiple Speech Streams
To answer RQ2, we trained PSLM variants with two (-2x) or three (-3x) speech streams [^5] PSLM-3x was trained with a batch size of 4 due to the increased number of parameters.. PSLM-2x achieved comparable scores to PSLM, whereas PSLM-3x demonstrated significant degradation. From these results, we conclude that speech tokens can be decoded in up to two streams without quality degradation (RQ2). An ablation study can be found in Appendix D.
4.2 Human Evaluation
Considering practical applicability to SQA, we manually evaluated three methods: CoM-SQ, CoM-ASR, and PSLM-ASR, which do not rely on gold TQ, along with Ground-Truth. Table 2 shows the results. The text response naturalness of PSLM-ASR was comparable to CoM-ASR and higher than CoM-SQ, which is consistent with the automatic evaluation results. For speech naturalness, all methods achieved higher scores than Ground-Truth. This result can be attributed to two reasons: (1) SA of Ground-Truth are synthetic speech, which may include errors in pronunciation, intonation, and pauses, and (2) SA of Ground-Truth are typically longer than those of other methods, incurring that one or two unnatural parts lowered the entire score. Nevertheless, we confirmed that our approach can generate natural and faithful speech responses. For response speed evaluation, PSLM-ASR achieved a significantly higher score than CoM-ASR and CoM-SQ. This finding verifies that the proposed method reduces latency both numerically and perceptibly. Detailed analysis can be found in Appendix E.2.
Table 2: Human evaluation results.
[…]
Ground Truth 4.08$\pm$0.26 3.74$\pm$0.19 4.73$\pm$0.11
CoM-SQ 2.44$\pm$0.29 4.04$\pm$0.20 4.07$\pm$0.23
CoM-ASR 2.90$\pm$0.30 3.94$\pm$0.20 4.17$\pm$0.22
PSLM-ASR 3.08$\pm$0.27 4.08$\pm$0.20 4.57$\pm$0.13
5 Conclusion
In this study, we proposed the Parallel Speech Language Model (PSLM), an LLM capable of generating text and speech tokens in parallel with multiple input-output streams, and investigated its impact on response quality and overall latency. The experimental evaluations on spoken question answering demonstrated that the proposed method significantly reduces latency compared to existing methods while maintaining response quality. Future work includes verifying the effectiveness of the proposed method on larger datasets and real speech data. Additionally, extending the proposed method to multi-turn dialogues is an important research direction.
6 Limitations
We recognize several limitations of this study. First, PSLM sacrifices ASR capability for faster response, requiring an external ASR module to serve as a spoken dialogue system. Enabling ASR with the PSLM architecture can be an interesting research direction. Second, we used single-speaker synthetic speech for SQ and SA, which lacks diversity in several aspects of speech such as accent, rhythm, emotion, and timbre. Practical applications may require to accept voices of arbitrary speakers, which we will address in future work. Finally, multi-turn dialogue settings were not investigated in our experiments. While SpeechGPT (Zhang et al., 2023) was not applied to multi-turn dialogue due to sequence length limitations, our models with multiple speech streams have the potential to perform multi-turn dialogue.
Appendix A Sequence Length Distributions
We calculated the sequence length distributions of SQ, TQ, TA, and SA in the training set. The results are listed in Table 3. On average, CoM prompting requires to generate $36.5\ (\textrm{TQ})+33.8\ (\textrm{TA})\approx 70$ text tokens before generating SA. Eliminating the need for generating these tokens can greatly reduce overall latency. In addition, speech tokens are more than 11 times longer than text tokens, highlighting the need for efficient generation of speech tokens.
Appendix B Implementation Details of HiFi-GAN
The HiFi-GAN generator comprises convolution layers. Therefore, a waveform fragment corresponding to the $i$-th token depends only on tokens with indices $[i-\lfloor R/2\rfloor,i+\lfloor R/2\rfloor]$. This allows waveform generation to start before the entire SA is generated. As described in Figure 2, HiFi-GAN first generates a waveform fragment once the LM generates $N_{\textrm{offset}}=\lfloor R/2\rfloor+1$ tokens, then generates subsequent fragments by shifting input tokens one by one.
In our experiments, we trained HiFi-GAN to generate 24 kHz waveform from 50Hz tokens, which results in $R=26$. Following Polyak et al. (2021), we embedded input speech tokens into 256-dimensional features and fed them to HiFi-GAN. We modified the upsampling rates to $[8,6,5,2]$, the number of total iterations to 300k, and kept the other configuration the same as the original work (Kong et al., 2020).
Appendix C ChatGPT Evaluation Prompt
We used the prompt in Figure 3 for ChatGPT-based evaluation. The original prompt was written in Japanese, but a translated version is presented here.
Appendix D Ablation Study
We trained three PSLM variants, one from scratch (-no-pretrain), one without TQ (-no-TQ), and one without SQ (-no-SQ). In addition, we trained PSLM-2x and PSLM-3x without weighted loss (-no-WL). Table 4 shows the automatic evaluation results. PSLM-no-pretrain exhibited significant degradation in all metrics, indicating the necessity of pretrained LM’s text capability. PSLM-no-TQ also showed large degradation, highlighting the importance of TQ in response quality. In contrast, PSLM-no-SQ achieved comparable scores to PSLM. This result implies that the speech-specific information such as intonation, rhythm, and emotion is not essential in the current SQA task due to the use of synthetic speech. We also found that PSLM-2x-no-WL achieved almost comparable scores to PSLM, whereas PSLM-3x-no-WL showed significant degradation. From these results, we conclude that the weighted loss is especially effective as the number of speech streams increases.
Table 3: Sequence length distributions in the training set (in tokens).
SQ TQ TA SA
Mean 406.6 36.5 33.8 386.5
Min 34 2 1 27
25% 214 19 15 179
50% 354 32 29 340
75% 577 51 50 563
Max 1861 148 147 1697
Figure 2: Streaming inference using HiFi-GAN with receptive field size $R=5$ and SA length $N_{\textrm{SA}}=6$. Waveform generation begins once $N_{\textrm{offset}}=\lfloor R/2\rfloor+1=3$ tokens are generated. Text tokens are omitted.
Appendix E Latency Analysis
E.1 Configuration for Latency Simulation
For latency calculation, we set $D_{\textrm{s2t}}=0.05$, $D_{\textrm{SQ}}=0.05$, $D_{\textrm{ASR}}=0.2$, and $D_{\textrm{t2s}}=0.01$ based on measurements on a single NVIDIA A100 GPU. The actual TPS varies with resources and optimization; 70 TPS was achieved with vLLM (Kwon et al., 2023) optimization, and 25 TPS without it. Meanwhile, to enable streaming inference with HiFi-GAN, LMs need to generate 50 speech tokens per second. Thus, we set $P$ to 50 for Table 1 and 50 or 100 in the next section.
E.2 Detailed Latency Analysis
We investigated the overall latency for different TA lengths. Figure 4 shows the results. Due to the need for prior generation of TA, the latency of CoM-SQ and CoM-ASR increases linearly as TA length increases. In contrast, the latency of PSLM-ASR is constant because Equation 3 does not include $N_{\textrm{TA}}$, and PSLM-2x-ASR further reduces the latency. The gap between CoM-based and PSLM-based systems is remarkable when generating long TA, highlighting the effectiveness of generating text and speech tokens in parallel.
Figure 3: A prompt for ChatGPT evaluation.
Table 4: Ablation study. The suffix no-WL denotes weighted loss was not applied.
[…]
PSLM SQ, TQ (Gold) TA, SA 3.50$\pm$0.08 3.22$\pm$0.09 5.05 5.25
PSLM-2x SQ, TQ (Gold) TA, SA 3.50$\pm$0.08 3.20$\pm$0.09 4.29 6.39
PSLM-3x SQ, TQ (Gold) TA, SA 3.28$\pm$0.10 2.99$\pm$0.10 7.07 6.09
PSLM-no-pretrain SQ, TQ (Gold) TA, SA 2.22$\pm$0.07 2.12$\pm$0.07 18.18 10.13
PSLM-no-TQ SQ TA, SA 2.34$\pm$0.09 2.19$\pm$0.09 8.84 6.38
PSLM-no-SQ TQ (Gold) TA, SA 3.54$\pm$0.08 3.17$\pm$0.09 6.31 8.99
PSLM-2x-no-WL SQ, TQ (Gold) TA, SA 3.42$\pm$0.08 3.17$\pm$0.08 8.84 4.99
PSLM-3x-no-WL SQ, TQ (Gold) TA, SA 2.67$\pm$0.10 2.46$\pm$0.10 11.36 6.94
Figure 4: Latency vs. TA length for different methods and token generation speed. PSLM-2x-ASR (50 TPS) is omitted because its latency is identical to PSLM-ASR (100 TPS).